OpenStudy (anonymous):

Express $63 as a percentage of $210 Express $63 as a percentage of $210 @Mathematics

OpenStudy (anonymous):

Let x be the percentage 210 x = 63 x = 63/210 = 0.3 30 %

OpenStudy (anonymous):

63/210 = 0.3 0.3 * 100 (percentage means, out of 100) so =30 %